Equvalate the Tan{(-1)^n π/4}=? Where n is integer.

Grade:11

1 Answers

Vikrant Kumar
7 years ago
if n is even
for eg: let n=2 then
Tan{(-1)^2 π/4}= Tan{π/4}=1
and for n being odd
let n=1
then, Tan{-π/4}= -Tan{π/4}= -1Tan{(-1)^1 π/4}=
this is a general solution for all n belonging to even and odd numbers